OpenStudy (anonymous):

Which of the systems below is equivalent to the given system? 0.01x - 0.05y = 0.1 0.4x + 0.3y = 0.9

OpenStudy (anonymous):

10x - 5y = 10 and 4x + 3y = 9 x - 5y = 10 and 4x + 3y = 9 x - 5y = 1 and 4x + 3y = 9

OpenStudy (anonymous):

10x-5y=10 and 4x+3y=9

OpenStudy (callisto):

For the equation 0.01x - 0.05y = 0.1 Multiply both sides by 100, what do you get?

OpenStudy (anonymous):

x-5y=1

OpenStudy (callisto):

Left side => correct Right side => not correct 0.1 x 100 = ?

OpenStudy (anonymous):

=10 whoops typo

OpenStudy (callisto):

So, you'll get the first equation now. 0.4x + 0.3y = 0.9 Next, multiply both sides by 10, what do you get?

OpenStudy (anonymous):

4x+3y=9

OpenStudy (anonymous):

so u get x - 5y = 10 and 4x + 3y = 9

OpenStudy (callisto):

Yes